# Two dimensional dimers beyond planarity

### v2

@@ Abstract
We consider a generalisation of the double dimer model which includes several models of interest, such as the monomer double dimer model, the dimer model, the Spin O(N) model, and it is related to the loop O(N) model. We prove that on two-dimension like graphs (such as slabs), both the correlation function and the probability that a loop visits two vertices converge to zero as the distance between such vertices gets large. Our analysis is by introducing a new (complex) spin representation for all models belonging to this class, and by deriving a new proof of the Mermin-Wagner theorem which does not require the positivity of the Gibbs measure. Even for the well studied dimer and double dimer model our results are new since - not relying on solvability and Kasteleyn’s theorem - they hold beyond the framework of planar graphs.
@@ 1 Introduction
The dimer model is the study of the set of perfect matchings of a graph. It attracts interest from a wide range of perspectives, including probability, combinatorics, statistical mechanics, graph theory. The double dimer model consists in the superposition of two perfect matchings of a graph drawn independently at random with uniform distribution, its realisations can be viewed as collections of disjoint self-avoiding loops.
@@
The rigorous study of such models achieved a breakthrough with the seminal works of Kasteleyn [15], Temperley and Fisher [31], who proved a theorem (known as Kasteleyn’s theorem) allowing an exact computation of the number of dimer configurations on planar graphs. Understanding the properties of perfect matchings beyond the framework of Kasteleyn’s theorem is much harder and of great interest. New results in this direction have recently been achieved in the study of the dimer problem in $\mathbb{Z}^{d}$ for $d\geq 3$ [6, 28, 30, 27] and on weakly planar graphs [19].
The present paper considers the dimer model, the double dimer model and a generalisation of such models on two dimensional (not necessarily planar) graphs. We will discuss the most general setting in Theorem 7.2 and Section 7.1 below, will now focus our discussion on the monomer double dimer model, a special case of our general framework which is of central interest. The configurations of the monomer double dimer model are obtained by selecting a random set of vertices (monomers) and then drawing at random two independent sub-graphs such that the degree of each monomer is zero and the degree of the remaining vertices is precisely one. The realisations of such a model can be viewed as disjoint collections of self-avoiding loops (see the left of Figure 1.1 for a configuration with four monomers as an example, where the edges of the two sub-graphs are coloured by blue and red respectively). The model depends on a parameter, the monomer activity, which rewards the total number of monomers. When the monomer activity is zero no monomer is present and the model reduces to the double dimer model.
Figure 1.1: Left: A configuration of the monomer double dimer model in a box of $\mathbb{Z}^{2}$. Right: A dimer configuration on the duplicated graph. The red, orange and blue dimers represent the vertical dimers and the dimers of the lower and upper box, respectively. Projecting the upper and lower box onto the same box leads to the configuration on the left.
If the monomer activity is positive, it is not necessarily the case that each vertex is touched by a loop. Alternatively, the configurations of this model on a graph $G$ can be viewed as a perfect matching of a graph corresponding to two copies of $G$, with each vertex of the first copy being connected by an edge to the corresponding vertex of the second copy. The monomers in the first formulation then correspond to dimers on the edges connecting the two copies in the second formulation, as it is represented in Figure 1.1.
The presence a positive density of monomers makes the model computationally intractable [11] and Kasteleyn’s theorem inapplicable, understanding simple properties of this basic model is then of great interest.
We state our results on two dimensional slabs with arbitrary width (the width may even ‘slowly’ grow to infinity with the number of vertices), the two-dimensional planar torus is then a special case of our theorems.
Our first main theorem (Theorem 1.1 below) provides an upper bound for the typical length of the loops in the monomer double dimer model. Our theorem states that, for each choice of the monomer activity, the expected loop length divided by the number of vertices in the graph converges to zero as the number of vertices in the graph tend to infinity. Our result then implies that in two dimensional graphs the probability that two vertices are connected by the same loop decays to zero with their distance, contrary to dimensions three and higher [27].
Our second main theorem (Theorem 1.2 below) involves monomer-monomer correlations in the dimer model, namely the ratio between the number of perfect matchings of a graph from which two vertices (and all the edges connected to them) have been removed and the number of perfect matchings of the graph. We show that the Cesaro sum of monomer-momomer correlations converges to zero as the number of vertices in the graph tend to infinity.
Notice that our first main theorem is known in the special case of zero monomer activity (the double dimer model) on planar graphs [9, 22] and our second main theorem is known for the dimer model on planar graphs [8]. The novelty of our work is to prove such results beyond the framework of Kasteleyn’s theorem, e.g. on two dimensional non-planar graphs and/or in the presence of a positive density of monomers. Indeed, even on planar graphs, the dimer model with a positive density of monomers is generally not solvable. Only if the monomers are present on the boundary of a bounded domain Kasteleyn’s theorem is applicable, since in this case one can map the system into a dimer model in an enlarged graph (see [2, 18] for recent works).
Our results hold in even more general frameworks. For example, they hold for the multiple dimer model, with dimers having $2N$ possible ‘colours’, such that the number of dimers of even colour $i$ touching any vertex is arbitrary and equals the number dimers of colour $i+1$ touching the same vertex. The monomer double dimer model then corresponds to the special case $N=1$, with the number of dimers of colour $1$ touching each vertex being equal to the number of dimers of colour $2$ touching the same vertex and this number being either 00 or $1$. We treat such generalisations in Theorem 7.2 and in Section 7.1 below. Such a general framework also includes a model which is closely related to the loop O(N) model (when $N$ is even), which has been mostly studied on the hexagonal lattice (see e.g. [10, 26] and references therein). The (minor) difference between our model and the loop O(N) model on the hexagonal lattice is that in the latter one no loop of length two is present. Our techniques are complementary to those in [10], since they provide a lower bound for the probability of observing a long loop, and, contrary to [10], our method can be applied only for integer values of $N$, but for any value of the inverse temperature parameter and even on graphs which are not necessarily hexagonal (and which are symmetric enough to employ the reflection positivity method, for example periodic boundary conditions are a necessary ingredient).
A key ingredient of our proofs is a spin representation relating connection probabilities and monomer correlations in the monomer double dimer model to spin-spin correlations. Our spin representation presents a complex duality relation between the two spin components which plays a crucial role in our analysis and makes the representation invariant with respect to global spin rotations. The main obstacle in using such representation is that, since the measure is non positive (and complex), one cannot employ the classical strategies [7, 14, 17, 25] based on the spin-wave deformation, or directly apply the older approach of [13]. Indeed, these approaches use the Cauchy-Schwarz inequality in an essential way, which does not produce useful bounds for non-positive measures. We overcome these obstacles introducing a new proof of a Mermin-Wagner theorem for such complex measures. We rely on the reflection positivity property of such models, which allows us to derive a weak form of a Cauchy-Schwarz inequality for a restricted (but useful) class of functions (see Theorem 6.3 below). This inequality is used to show that the magnetisation vanishes if we take the limit of small external fields appropriately. A perturbative analysis allows us to show that a vanishing magnetisation implies that the Cesaro sum of two point functions is small (see Proposition 5.4 below). Our general strategy for proving the Mermin Wagner theorem is new and can then be applied to any spin system with a non necessarily positive (but reflection positive) measure.
Parallel to our work, a related complex spin representation has been introduced in [3] for the derivation of infrared bounds. The results which are presented in [3] have been obtained independently from us.
Main results.
@@
Consider a finite undirected bipartite graph $G=(V,E)$. A dimer configuration is a spanning sub-graph of $G$ such that every vertex has degree one. We let $\mathcal{D}_{G}$ be the set of dimer configurations in $G$. Given a set $A\subset V$, we let $G_{A}$ be the subgraph of $G$ with vertex set $V\setminus A$ and with edge set consisting of all the edges in $E$ which do not touch any vertex in $A$. We let $\mathcal{D}_{G}(A)$ be the set of dimer configurations in $G_{A}$.
The configuration space of the monomer double dimer model is denoted by $\Omega$ and it corresponds to the set of triplets $\omega=(M,d_{1},d_{2})$ such that $M\subset V$ and $(d_{1},d_{2})\in\mathcal{D}_{G}(M)\times\mathcal{D}_{G}(M)$. We refer to the first element of the triplet $\omega$ as a set of monomers, and to the second and third element as set of dimers. For example, a configuration with four monomers in the box of $\mathbb{Z}^{2}$ can be viewed in the left of Figure 1.1. As one can see from the figure, any configuration $(M,d_{1},d_{2})\in\Omega$ can be viewed as a collection of disjoint self-avoiding loops, with loops possibly having length two and corresponding to the superposition of two dimers on the same edge. We let $\mathcal{M}:\Omega\mapsto V$ be the random variable corresponding to the set of monomers, i.e, for each $\omega=(M,d_{1},d_{2})\in\Omega$ $\mathcal{M}(\omega):=M$. We define a probability measure on $\Omega$,
$\forall\omega\in\Omega\quad\mathbb{P}_{G,N,\rho}(\omega):=\frac{\rho^{|\mathcal{M}(\omega)|}\,N^{\mathcal{L}(\omega)}}{\mathbb{Z}_{G,N,\rho}},$ (1.1)
where $\rho,N\geq 0$ are some parameters, $\mathcal{L}(\omega)$ is the number of loops in $\omega$ and $\mathbb{Z}_{G,N,\rho}$ is a normalisation constant. We refer to $\rho$ as the monomer activity and use $\mathbb{E}_{G,N,\rho}$ to denote the expectation with respect to (1.1). The monomer double dimer model illustrated in Figure 1.1 corresponds to the special case of $N=1$.
We define $L_{x}={L}_{x}(\omega)$ as the loop which touches the vertex $x\in V$ in the configuration $\omega\in\Omega$ and set $L_{x}(\omega)=\emptyset$ if the configuration $\omega$ has a monomer at $x$ Moreover, we denote by $|{L}_{x}|=|{L}_{x}|(\omega)$ the length of the loop $L_{x}(\omega)$, namely the number of vertices which are touched by such a loop, and set $|{L}_{x}|(\omega)=0$ if $\omega$ has a monomer at $x$.
@@
We let $\mathbb{S}_{L,K}$ denote the two dimensional slab torus, which we identify with the set $\big{(}(-\frac{L}{2},\frac{L}{2}]\cap\mathbb{Z}\big{)}\times\big{(}(-\frac{L}{2},\frac{L}{2}]\cap\mathbb{Z}\big{)}\times\big{(}(-\frac{K}{2},\frac{K}{2}]\cap\mathbb{Z}\big{)}$. In the special case $K=1$, $\mathbb{S}_{L,1}$ corresponds to the planar torus, while $\mathbb{S}_{1,1}$ is a single point. We use $o\in\mathbb{S}_{L,K}$ to denote the origin. In (1.1) and in the corresponding expectation we use the subscript ${}_{L,K}$ rather than ${}_{\mathbb{S}_{L,K}}$ in order to simplify the notation. Our theorem states that the expected loop length divided by the volume converges to zero in the limit of large slabs.
Theorem 1.1.
Suppose that $\rho>0$ and $N\in\mathbb{N}$ or that $\rho=0$ and $N=1$. There exists $c=c(\rho,N)<\infty$ such that, for any $L,K\in 2\mathbb{N}\cup\{1\}$ satisfying $K\leq\sqrt{\log(L)}$,
$\frac{\mathbb{E}_{L,K,N,\rho}\big{(}|L_{o}|\big{)}}{|\mathbb{S}_{L,K}|}\leq c\sqrt{\frac{K}{\log L}}.$ (1.2)
Since $\mathbb{E}_{L,K,N,\rho}\big{(}|L_{o}|\big{)}=\sum_{x\in\mathbb{S}_{L,K}}\mathbb{P}_{L,K,N,\rho}(o\leftrightarrow x)$, our theorem implies that the probability that two distant vertices are visited by the same loop is not uniformly positive, contrary to dimension three and higher and small monomer activity [28].
Our previous theorem can also be stated for other quantities of interest in the monomer double dimer model with arbitrary monomer activity, such as the Cesaro sum of the two-point function $\mathcal{G}^{1}_{G}(x,y)$ which is introduced in equation (3.5) below. In the special case of zero monomer activity such a two-point function reduces to the monomer-monomer correlation in the dimer model,
$\forall x,y\in V\quad\mathcal{C}_{G}(x,y):=\frac{|\mathcal{D}_{G}(\{x,y\})|}{|\mathcal{D}_{G}|},$ (1.3)
@@
a quantity of central interest in the study of this model (if $G=\mathbb{S}_{L,K}$, we use the subscript ${}_{L,K}$ rather than ${}_{\mathbb{S}_{L,K}}$, in order to simplify the notation). Our third main theorem provides a upper bound on the Cesaro sum of such a function on two dimensional slabs.

### v3

# Mermin–Wagner theorem

@@ Abstract
We study a generalisation of the double-dimer model that encompasses several models of interest, including the monomer double-dimer model, spatial random permutations, the dimer model, and the spin $O(N)$ model, and which is also related to the loop $O(N)$ model. We show that on two-dimensional-like graphs (such as slabs), both the correlation function and the probability that a loop visits two vertices decay to zero as the distance between the vertices diverges. Our approach is based on the introduction of a new complex spin representation for all models in this class, together with a new proof of the Mermin–Wagner theorem that does not require positivity of the Gibbs measure. Even for the well-studied dimer and double-dimer models our results are new: since they do not rely on exact solvability or Kasteleyn’s theorem, they apply beyond the planar-graph setting.
@@ 1 Introduction
We study a general class of statistical mechanics models, naturally described in terms of random subgraphs or collections of interacting random walks on two-dimensional (not necessarily planar) graphs. This framework encompasses three paradigmatic examples: the dimer model, spatial random permutations, and the monomer double-dimer model. The main result is a Mermin–Wagner theorem for the whole class, from which the decay of correlations and absence of long-range order or of macroscopic loops on two-dimensional graphs follow, and which is new for all these models.
1.1 Main results
In what follows, we provide a brief description of these models and discuss the novelty of our results in each case.
The dimer model.
The dimer model is the study of the set of perfect matchings of a graph. It has attracted interest from a wide range of perspectives, including probability, combinatorics, statistical mechanics, and graph theory. Configurations of the model are subgraphs of a fixed (bipartite) graph in which every vertex has degree exactly one, and they are sampled uniformly at random.
@@ 1.1 Main results
A major breakthrough in the rigorous study of such models was achieved through the seminal works of Kasteleyn [28], Temperley and Fisher [52], who established a theorem (now known as Kasteleyn’s theorem) that allows an exact computation of the number of dimer configurations on planar graphs.
Understanding the properties of the dimer model beyond the planar setting is significantly more challenging and of considerable interest. In this case the model is no longer integrable and is even computationally intractable [27]. Nevertheless, important progress has recently been made in the study of the dimer problem on $\mathbb{Z}^{d}$ for $d\geq 3$ [12, 31, 46, 47, 50] and on weakly planar graphs [22].
In this work, we investigate the dimer model on two-dimensional not necessarily planar graphs and address the fundamental problem of correlation decay. More precisely, we study monomer–monomer correlations, a central quantity in the rigorous analysis of the model [18], and establish a Mermin–Wagner theorem, which implies decay of correlations and, consequently, the absence of long-range order—in sharp contrast with the behavior in dimensions $d>2$ [50]. Our main result for the dimer model is stated in Theorem 1.1 below. In the planar setting the decay of monomer-correlations was conjectured in [18] and established in [15] using Kasteleyn’s theorem and the study of Cauchy-Riemann operators.
The monomer double-dimer model.
Figure 1.1: Left: A configuration of the monomer double-dimer model in a box of $\mathbb{Z}^{2}$. Center: A configuration of the spatial permutation model in a box of $\mathbb{Z}^{2}$, where fixed points of the permutation are represented by isolated vertices. This configuration corresponds to the one on the left. Right: A configuration of the monomer double-dimer model in $\Omega_{o,x}$ when $x$ and $o$ have odd distance. Since $o$ and $x$ are the only vertices incident to exactly one dimer, while all other vertices are incident to either none or two dimers, the existence of a self-avoiding path with $o$ and $x$ as endpoints follows.
The double-dimer model arises by superimposing two independent dimer configurations on the same graph. The resulting random structure consists of loops and doubled edges, and provides a natural extension of the classical dimer model. Beyond its combinatorial appeal, the model has attracted considerable attention due to its deep connections with conformal invariance and loop ensembles. In particular, it is conjectured that that loops in the planar double-dimer model converge in the scaling limit to Conformal Loop Ensembles $\mathrm{CLE}(4)$, Kenyon proved that the scaling limit is conformal invariant [30], while Dubédat established relations to isomonodromic deformations and conformal field theory [16]. These results place the double-dimer model at the crossroads of probability, statistical mechanics, and complex analysis, and motivate the study of its behavior beyond the planar setting [22, 47].
On planar graphs, the double-dimer model is integrable and therefore amenable to precise analysis. The picture changes dramatically once monomers are allowed. In the monomer double-dimer model, one samples a random subset of vertices (“monomers”) that are removed together with all incident edges, and then a double-dimer configuration on the induced complementary subgraph (see Figure 1.1, left). A parameter—the monomer activity—controls the monomer density. In the presence of monomers, Kasteleyn’s theorem [28] no longer applies on planar or nonplanar graphs; the classical Kasteleyn framework only covers the special case when all monomer insertions lie on the boundary (see, e.g., [10]) and does not treat the general monomer double-dimer setting. Consequently, this model remains poorly understood and presents a challenging and compelling direction for future research.
Our main result, Theorem 1.2, establishes the absence of long-range order in the monomer double-dimer model on two-dimensional graphs, for every value of the monomer activity. More precisely, we prove decay to zero of the two-point function and the absence of macroscopic loops, namely loops whose size is proportional to the size of the box. This stands in sharp contrast with the behavior in dimension $d>2$, where long-range order is known to occur and macroscopic loops are present with uniformly positive probability [50, 46]. Remarkably, our result is new even for the well-studied double-dimer model, that is, the special case of zero monomer activity, since it applies to non-planar graphs where integrability is no longer available.
Spatial random permutations.
Another object of central interest is that of spatial random permutations on a lattice. The configuration space is given by the set of permutations of the vertices of a finite box, with the restriction that each vertex is mapped either to itself or to one of its nearest neighbours. Each configuration can be represented as a collection of mutually disjoint directed self-avoiding cycles of even length, together with vertices that are mapped to themselves (the fixed points of the permutation), see Figure 1.1 for an illustration. There is a bijection between configurations of spatial random permutations on a lattice and those of the monomer double-dimer model; the article is written for the latter, but the results apply naturally to spatial permutations as well thanks to this correspondence (see Figure 1.1).
Random permutations are classical objects of study in probability theory and combinatorics. The spatial version has been proposed as a toy model for the Bose gas [4, 5, 13, 24, 49, 54]. In these systems loops interact through a potential depending on their total number and of their length. Such interactions make the model both interesting and challenging; importantly, however, they do not depend on the mutual spatial displacement of the loops, in contrast with the models studied here. The presence of spatial interactions, arising from the fact that loops are self-avoiding and mutually disjoint, makes the analysis of the model considerably more challenging and conceptually closer to that of the Bose gas.
The study of spatial random permutations in two dimensions has previously been addressed numerically in [6], where a variant of the model was studied: all permutations of the vertices of the box are allowed, but the measure is weighted with Gaussian factors penalizing jumps according to their length, thus effectively allowing only jumps of finite size, independently of the size of the box, just as in the version considered here. Based on numerical evidence, the author conjectures the existence of a Kosterlitz–Thouless transition, and hence the absence of long-range order and of macroscopic cycles at any temperature.
In this article we present the first rigorous study of spatial random permutations on two-dimensional graphs, we prove a Mermin–Wagner theorem for this model, and deduce the absence of long-range order and of macroscopic cycles at any temperature. This is in sharp contrast with the situation in dimension $d>2$, where long-range order has been proved to occur at sufficiently low temperature [50, 47] and where the existence of macroscopic cycles has been established in the special case of fully packed loops, that is, when fixed points of the permutation are suppressed.
Generalisations of our results.
Our results extend to more general frameworks of random loops.
A first generalisation consists in considering a model with dimers of two types, say red and blue, where the number of blue dimers incident to each vertex equals the number of red dimers. Unlike in the monomer double-dimer model, this number is not restricted to 0 or $1$, but can be any non-negative integer. By introducing additional “matching” variables at each site, which match every blue dimer with a corresponding red dimer, configurations can once again be viewed as collections of interacting loops. We refer to this model as multi-occupancy double-dimer model. All our results extend to this more general setting, and the article is in fact written in this framework.
A further generalisation covered by our results allows not just two dimer types but $2N$ types (labelled $1,\ldots,2N$) with $N\in\mathbb{N}$. For each odd $i\in\{1,3,\ldots,2N-1\}$ we impose the local constraint that, at every vertex, the number of incident dimers of type $i$ equals the number of incident dimers of type $i+1$. Introducing auxiliary pairing variables that match type $i$ only with type $i+1$, any configuration decomposes into a collection of loops, each alternating between colours $i$ and $i+1$ for some odd $i$. This framework also encompasses a model closely related to the loop $O(N)$ model (for even $N$), which has been studied primarily on the hexagonal lattice (see, e.g., [37, 42]). The only (minor) distinction is that in the hexagonal loop $O(N)$ model loops of length two are excluded. Our techniques are complementary to those developed in [37]: while the latter provide lower bounds on the probability of long loops, our methods yield upper bounds. Moreover, unlike [37], our approach applies only to integer values of $N$, but it works for any value of the inverse temperature parameter and extends beyond the hexagonal lattice, as long as the graph is symmetric enough to allow for reflection positivity (for instance, periodic boundary conditions are required). We address these generalisations in Section 8.1.
1.2 Discussion
Mermin–Wagner theorem and spin representation.
A key ingredient of our proofs is a spin representation that relates connection probabilities and monomer correlations in the monomer double-dimer model to spin–spin correlations. Our representation exhibits non-trivial complex dualities between the two spin components (see Remark 2.1 below), which on the one hand ensure reflection positivity of the measure and on the other hand make the model invariant under global spin rotations. This framework further allows one to perform a spin-wave approximation and to heuristically study dimer correlations in terms of a Gaussian free field or a Coulomb gas.
The main obstacle to turning such heuristics into rigorous results is that, since the measure is complex and nonpositive, the classical strategies based on spin-wave deformations [14, 20, 40, 38] cannot be employed. Nor can one directly apply the earlier approach of [39], which also relies on positivity assumptions.
To put this difficulty in perspective, recall that there are essentially two established approaches to the proof of the Mermin–Wagner theorem in the literature. One approach, due to Pfister [43] (see also the exposition in [42] and [55]), is based on entropy estimates and relies crucially on the positivity of the Gibbs measure (see also, [41] for an extension to degenerately positive Gibbs measures ). Alternatively, the original argument of Mermin and Wagner (see [40]) is built upon the Bogoliubov functional inequality. This method has been extended in several directions, e.g., to quantum spin systems [3] and to certain supersymmetric spin systems that admit a positive Gibbs measure representation [2]. In the latter case, the spins take values in a hyperbolic space, which can be mapped to a positive Gibbs measure, so that the classical proof can be adapted ((see also, the work of Kozma and Peled [32] and Sabot [48] for the polynomial decay for the vertex-reinforced jump process via similar spin representations). At the core of this line of reasoning lies a Bogoliubov-type inequality derived via a scalar or operator Cauchy–Schwarz inequality under positivity of the Gibbs (or quantum Gibbs) measure.
In contrast, our contribution is to provide a new proof of a Bogoliubov-type inequality, Theorem 6.3, based on the observation that the Fourier transform of an (almost) convex function on the lattice is (almost) nonnegative. This perspective does not require positivity of the measure and therefore opens the door to extending Bogoliubov-type inequalities (and, consequently, the Mermin–Wagner theorem) to more general settings such as reflection-positive measures. In this way, our approach both complements and extends the existing frameworks in a novel direction.
Our version of the Mermin–Wagner theorem is presented in Theorem 7.1. It states that, at any monomer activity, the magnetisation vanishes when the external field is taken to zero in a suitable way in two dimensions. A perturbative analysis then shows that vanishing magnetisation implies that the Cesàro sum of two-point functions is small (see Proposition 5.5), thus leading, via the Mermin–Wagner theorem, to the proof of our main results, Theorems 1.1 and 1.2.
Overview on other related models.
As explained above, one of the central objects of the present work is the monomer double-dimer model, that is, a configuration consisting of a set of monomers together with two independent dimer configurations on the complementary set of vertices. It is therefore natural to compare it with other classical models of statistical mechanics obtained by superimposing monomers and dimers.
A first example is the monomer–dimer model, defined as the superposition of a set of monomers and a single configuration of dimers on the complementary set. Configurations are sampled at random according to a measure depending on a parameter, the monomer activity, which controls the density of monomers. The model reduces to the pure dimer model when the monomer activity is set to zero.
This model has been studied in several works, starting from [26], where a Lee–Yang theorem was obtained for any value of the monomer density. A natural consequence of this theorem is that correlations decay exponentially fast at all strictly positive values of the monomer activity [45] (see also [8] for a probabilistic proof of this result). Hence, in contrast to the monomer double-dimer model — which is conjectured to exhibit a Kosterlitz–Thouless transition in two dimensions and is known to display long-range order at low monomer activity in three dimensions [47] — the monomer–dimer model does not exhibit any non-trivial phase transition as the monomer activity varies, in any dimension. An interesting open problem for this model is to characterise how the mass gap depends on the monomer activity; see [45, Remark 2.2] for a discussion.
Another natural variant is the double monomer–dimer model, obtained by superposing two independent samples of the monomer–dimer model on the same graph. Because the monomer locations in the two configurations need not coincide, the superposition is not merely a union of disjoint cycles and monomers. Instead, three types of structures appear: (i) monomers (vertices that carry a monomer in both configurations); (ii) closed cycles with edges alternating between the two samples; and (iii) open alternating paths whose two endpoints are monomers (one from each configuration). To the best of our knowledge, this model has not been analyzed in the literature. Nevertheless, the model admits a representation within our spin formalism at positive external field, with the field playing the role of the monomer activity (in contrast to the monomer double-dimer model, where this role is played by the temperature of the spin measure). We expect that, in this regime, no macroscopic paths occur in any dimension.
1.3 Definitions and main theorems
In this section we present our main results for the three paradigmatic examples of our general framework.
1.3.1 Dimer model
@@
Consider a finite undirected graph $G=(V,E)$. A dimer configuration is a spanning subgraph of $G$ such that every vertex has degree one. We let $\mathcal{D}_{G}$ be the set of dimer configurations in $G$ and suppose that $G$ is such that $\mathcal{D}_{G}$ is non-empty. Given a set $A\subset V$, we let $G_{A}$ be the subgraph of $G$ with vertex set $V\setminus A$ and with edge set consisting of all the edges in $E$ which do not touch any vertex in $A$. We let $\mathcal{D}_{G}(A)$ be the set of dimer configurations in $G_{A}$. We introduce the monomer-monomer correlation, namely the ratio between number of dimers covers of $G_{\{x,y\}}$ and of $G$.
$\forall x,y\in V\qquad\mathcal{C}_{G}(x,y):=\frac{|\mathcal{D}_{G}(\{x,y\})|}{|\mathcal{D}_{G}|}.$ (1.1)
@@
We let $\mathbb{S}_{L,K}$ denote the two-dimensional slab torus, which we identify with the set $\big((-\frac{L}{2},\frac{L}{2}]\cap\mathbb{Z}\big)\times\big((-\frac{L}{2},\frac{L}{2}]\cap\mathbb{Z}\big)\times\big((-\frac{K}{2},\frac{K}{2}]\cap\mathbb{Z}\big)$. When $K>1$, we then impose periodic boundary conditions in all three directions, each vertex has then $6$ neighbours. In the special case $K=1$ and $L>1$, $\mathbb{S}_{L,1}$ corresponds to the planar torus, each vertex has then $4$ neighbours, while $\mathbb{S}_{1,1}$ is a single point. We use $o\in\mathbb{S}_{L,K}$ to denote the origin. If $G=\mathbb{S}_{L,K}$, we use the subscript L,K rather than $_{\mathbb{S}_{L,K}}$, in order to simplify the notation, and we use the same convention for the other quantities defined below.
@@
Our main theorem on the dimer model provides a upper bound on the Cesaro sum of such a function on two-dimensional slabs.